Natural gas is transmitted by a spherical tank with an inner radius of ri and outer radius of ro. The tank material is stainless steel that has a thermal conductivity of k. The inner surface temperature is at Tw, and the outer surface dissipating heat by convection with a heat transfer coefficient h into the ambient air at temperature T∞.
a) Formulate the conduction heat transfer problem.
b) Develop an expression for temperature distribution T(r) within the tank material.
c) Find heat transfer between the tank and atmosphere.
Answer #1

(a (6) one dimensional sphemical anduction Yone Man (medt/2.1] h, TOO olem (54 dt/n] 50 m od d9/277=61 dT/top a/2 = -c/rtc2 T